# ValuesRAG: Enhancing Cultural Alignment Through Retrieval-Augmented Contextual Learning

### v1

@@ 4.1 Values and Demographic Summary Generation
To systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages. First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved (described in Section 3.1). In parallel, topic-based summaries and demographic summaries are generated separately. For topic-based summaries, values-related QA sets are used to produce summaries for each topic, while demographic summaries are generated using demographic-related QA sets:
@@ 4.2 Values Augmented Generation
Once the comprehensive summaries for each individual are generated in the previous step, we construct an augmented generation process for evaluating the validation question-answer data. For each validation question, we concatenate the corresponding individual’s values summary with the question itself, forming a context-rich input for the LLM:
@@ 4.3 Retrieval-based Values Alignment
To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval. The demographic data from both the train and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model. We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
Methods EVS GSS CGSS ISD LAPOP Africa Avg.
Baseline
Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
Role-Assignment + Few-Shot Learning 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
Ours
ValuesRAG($k=1$) 0.5960 0.7722* 0.5347* 0.6853* 0.4682 0.5904* 0.6078*
ValuesRAG($k=3$) 0.6020* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
ValuesRAG($k=5$) 0.6051* 0.7706* 0.5301* 0.7016* 0.5061* 0.5905* 0.6173*
ValuesRAG($k=10$) 0.6020* 0.7380 0.5317* 0.7014* 0.4686 0.5680* 0.6016*
Table 2: Accuracy scores for various methods compared with multiple baselines across six regional datasets, including the European Values Study (EVS), General Social Survey (GSS), Chinese General Social Survey (CGSS), Indian Survey Data (ISD), LAPOP, Afrobarometer, and the AmericasBarometer. $k$ indicates the number of summaries to be retrieved. Bold text indicates the best performance, underlined text indicates the second-best performance. * denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p$ <0.05) over all baseline model(s).
@@
Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score. The top-100 embeddings with the highest similarity scores are initially selected as candidates. Subsequently, we apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates. The reranking process evaluates the semantic relevance of the initial candidates based on a scoring function:
@@
$P_{\text{test}}$ is the final prompt, $D_{\text{test}}$ is the demographic information of the test individual, $\{R^{\prime}_{1},R^{\prime}_{2},\dots,R^{\prime}_{K}\}$ represents the top-$k$ reranked summaries, and $Q_{\text{test}}$ is the test question. $A_{\text{test}}$ denotes the generated answer for the test question, and $f_{\text{gen}}$ represents the generation function that leverages the structured input prompt.
This retrieval-based approach, followed by reranking, enhances reasoning by explicitly guiding the LLM to critically evaluate which retrieved values best align with the test individual’s demographic characteristics. The final prompts are then used to generate answers following the chain-of-thought prompting strategy, ensuring that the responses are contextually coherent and culturally aligned with the test individual’s profile.
@@ 5.1 Setup
Models Used
We employ GPT-4o-mini Achiam et al. (2023), a smaller version of GPT-4o, for our generation tasks, with the temperature parameter set to $0.7$ for balancing coherence and creativity. For retrieval task, we used the E5 (base) model Wang et al. (2022) [^2] https://huggingface.co/intfloat/e5-base to generate embeddings and retrieve the top-100 most relevant summaries of values by calculating cosine similarity. Additionally, we utilized the GTE-multilingual-reranker-base model Zhang et al. (2024) [^3] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base for reranking. This reranking model refines the selection process and ensures that the top-k most relevant summaries are selected from the initial top-100 retrievals.
Baseline Methods and Implementation
Our baseline methods include: (1) Zero-shot Inference, (2) the role-assignment-only approach Tao et al. (2024), (3) a few-shot learning method Choenni and Shutova (2024), and (4) a hybrid method that combines both (1) and (2). Specifically, for the role-assignment baseline, we use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data. For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts. The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set.
Evaluation Method
We used accuracy as the primary evaluation metric, converting multi-choice responses into binary form for consistency and simplicity by dividing responses into two distinct groups based on contrasting answer patterns, which effectively captures agreement or disagreement and aligns with the structure of values-related questions.
@@ 5.2 Experimental Analysis
In our experiments, we compare our method ValuesRAG with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2. We found that the role-assignment method generally surpasses both zero-shot and few-shot approaches. By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance. In contrast, few-shot learning, which depends on a small set of examples, often faces challenges in generalizing to unseen scenarios, as different values dimensions are not strongly related in many cases. However, while role-assignment offers clearer cultural context, it may oversimplify complex personal traits by imposing rigid, stereotype-like constraints on agent behavior. The hybrid method seeks to address these limitations by combining role assignment for consistency with few-shot prompts to introduce variety, but still falls short in capturing the full complexity of cultural values.
@@
In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific demographic and cultural data for each agent. Specifically, ValuesRAG (k=$3$) achieves the best results, with ValuesRAG (k=$5$) following closely behind. Although using $1$ or $10$ retrieved summaries also achieved better performance than other baselines, retrieving $3$ and $5$ summaries provides a more balanced trade-off between retrieval diversity and contextual relevance. This retrieval-augmented framework enables the model to dynamically incorporate richer, more nuanced information about each agent, avoiding the limitations of rigid demographic labels or small sample prompts. As a result, our model captures the interplay between individual beliefs, social contexts, and cultural norms more effectively. Evaluations across diverse test datasets demonstrate that ValuesRAG consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
5.3 Ablation Study: Impact of Values-Only Generation
@@
To validate the robustness of ValuesRAG, we perform an ablation study using only value-augmented generation, thereby isolating the impact of demographic summaries on the model’s performance. Specifically, we used the WVS validation set—separated from the training data as outlined in Section 3—to evaluate the models. Table 3 presents a comparison of our method, using only summaries of values, against four baseline methods. The results demonstrate that ValuesRAG consistently outperforms the baselines across all datasets, achieving the highest accuracy even relying solely on summaries of values.
Methods Accuracy
Zero-Shot 0.6176
Role-Assignment 0.6747
Few-Shot Learning 0.6359
Role-Assignment + Few-Shot Learning 0.6670
Values Augmented Generation-Only (Ours) 0.6894
@@
This result confirms the effectiveness and robustness of the values-augmented generation approach. ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses. Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined prompts or demographic anchors. The result showcases the framework’s scalability and adaptability, proving it can mitigate biases and generate culturally coherent outputs with minimal reliance on external context.
@@ 6 Conclusion
We present ValuesRAG, an innovative Retrieval-Augmented Generation framework that revolutionizes cultural values alignment through its comprehensive approach to values context analysis. While existing approaches rely on pre-trained knowledge with fixed demographic labels, often leading to oversimplified generalizations and biases, or few-shot prompting techniques that struggle with generalization due to limited correlations across value dimensions, ValuesRAG overcomes these weaknesses by dynamically retrieving and integrating granular demographic data with detailed individual-level cultural profiles. Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines across diverse regional QA datasets. By combining adaptive retrieval mechanisms with in-context learning and reranking strategies, ValuesRAG captures complex cultural nuances, reduces biases, and ensures contextually aligned responses. This structured design enables ValuesRAG to bridge the gap between generic LLM capabilities and the demands of culturally sensitive applications, providing a scalable and robust solution for real-world use cases.

### v2

@@ 4.1 Values and Demographic Summary Generation
systematically generate concise summaries of values and demographics for each individual, we process the dataset in three stages. First, the dataset is stratified by topics and split into train and validation sets, ensuring that the distribution of each topic is preserved (described in Section 3.1). In parallel, topic-based summaries and demographic summaries are generated separately. For topic-based summaries, values-related QA sets are used to produce summaries for each topic, while demographic summaries are generated using demographic-related QA sets:
@@ 4.2 Values Augmented Generation
Once the comprehensive summaries for each individual are generated in the previous step, we construct an augmented generation process for evaluating on the validation question-answer data. For each validation question, we concatenate the corresponding individual’s values summary with the question itself, forming a context-rich input for the LLM:
@@ 4.3 Retrieval-based Values Alignment
To dynamically assign relevant values to test individuals, we leverage demographic information as documents for retrieval. The demographic data from both the train (retrieval corpus) and test datasets are preprocessed into a structured context format, as described earlier, and embeddings are generated for each demographic context using a representation model. We first retrieve the top-100 most similar summaries of values for each test individual by computing the cosine similarity between the embeddings of the test and train demographics:
@@
Specifically, $E_{\text{test}}$ and $E_{\text{train}}$ represent the embeddings of the test and training (retrieval corpus) demographic contexts, respectively, and $\text{Sim}(\cdot,\cdot)$ denotes the cosine similarity score. The top-100 embeddings with the highest similarity scores are initially selected as candidates. Subsequently, we apply a reranking step to refine the selection and identify the most relevant summaries among the retrieved candidates. The reranking process evaluates the semantic relevance of the initial candidates based on a scoring function:
@@
Here, $P_{\text{test}}$ is the final prompt, $D_{\text{test}}$ is the demographic information of the test individual, $\{R^{\prime}_{1},R^{\prime}_{2},\dots,R^{\prime}_{K}\}$ represents the top-$k$ reranked summaries, and $Q_{\text{test}}$ is the test question. $A_{\text{test}}$ denotes the generated answer for the test question, and $f_{\text{gen}}$ represents the generation function.
This retrieval-based approach, followed by reranking, enhances reasoning by explicitly guiding the LLM to critically evaluate which retrieved values best align with the test individual’s demographic characteristics. The final prompts are then used to generate answers following the chain-of-thought prompting strategy, ensuring that the responses are contextually coherent and culturally aligned with the test individual’s profile. For the comprehensive implementation of the ValuesRAG, we provide the Algorithm 1, which systematically outlines the processes of values and demographic summary generation, values-augmented generation, and retrieval-based values alignment, as shown below:
Algorithm 1 Values Generation and Retrieval Process
1:Dataset $\mathcal{D}$ with topics and demographic QA pairs, Generative Model $f_{\text{gen}}$, Embedding Model $f_{\text{embed}}$, Reranking Model $f_{\text{rerank}}$, Retrieval Top-$K$
2:// Values and Demographic Summary Generation
3:for each individual $i\in\mathcal{D}$ do
4:     Generate topic-based values summaries $T_{i}^{j}$ for each topic $j$
5:     Generate demographic summaries $D_{i}$
6:     Combine $T_{i}^{j}$ into comprehensive values summary $S_{i}$
7:end for
8:// Values Augmented Generation
9:for each validation question $Q_{i}^{\text{val},k}$ of individual $i$ do
10:     Construct context $C_{i}=\text{concat}(S_{i},Q_{i}^{\text{val},k})$
11:     Augment context with demographic summary $D_{i}$
12:     Generate answer $A_{i}=f_{\text{gen}}(C_{i},D_{i})$
13:end for
14:// Retrieval-based Values Alignment
15:for each test individual $i$ do
16:     Compute embeddings $E_{\text{test}}=f_{\text{embed}}(D_{\text{test}})$
17:     Retrieve top-$100$ values summaries by similarity:
18:      $\text{Sim}(E_{\text{test}},E_{\text{train}})$
19:     Rerank top-$K$ summaries:
20:      $R^{\prime}_{k}=f_{\text{rerank}}(E_{\text{test}},E_{C_{j}})$
21:     Final prompt $P_{\text{test}}=(D_{\text{test}},R^{\prime}_{1},\dots,R^{\prime}_{K},Q_{\text{test}})$
22:     Generate answer $A_{\text{test}}=f_{\text{gen}}(P_{\text{test}})$
23:end for
Model Methods EVS GSS CGSS ISD LAPOP Africa Avg.
GPT-4o mini Zero-shot Inference 0.5566 0.6026 0.4019 0.6109 0.4195 0.3923 0.4973
Role-Assignment 0.5738 0.7564 0.4813 0.6164 0.4742 0.5563 0.5764
Few-Shot Learning 0.5271 0.6538 0.4631 0.5804 0.4220 0.4258 0.5120
Hybrid Method 0.5938 0.7292 0.5048 0.6330 0.4414 0.5305 0.5721
ValuesRAG 0.6020* 0.7781* 0.5387* 0.7001* 0.5030* 0.5953* 0.6195*
Gemini 1.5 Flash Zero-shot Inference 0.5419 0.6408 0.4502 0.6017 0.4149 0.4181 0.5113
Role-Assignment 0.5598 0.7493 0.4770 0.6048 0.4747 0.5262 0.5653
Few-Shot Learning 0.5225 0.6376 0.4559 0.5782 0.4194 0.4758 0.5149
Hybrid Method 0.5845 0.7193 0.5026 0.6253 0.4448 0.5166 0.5655
ValuesRAG 0.5869 0.7686* 0.5337* 0.6789* 0.4705 0.5473* 0.5977*
Table 2: Accuracy scores for various methods compared with multiple baselines across six regional datasets. $k$ indicates the number of summaries to be retrieved. Bold text indicates the best performance, underlined text indicates the second-best performance. * denotes significant improvements (paired $t$-test with Holm-Bonferroni correction, $p<0.05$) over all baseline model(s). denotes our proposed method.
Model Num(K) EVS GSS CGSS ISD LAPOP Africa Avg.
GPT-4o mini 1 0.5960 0.7722 0.5347 0.6853 0.4682 0.5905 0.6078
3 0.6021 0.7781 0.5387 0.7001 0.4686 0.5953 0.6138
5 0.6052 0.7706 0.5301 0.7016 0.5061 0.5905 0.6174
10 0.6020 0.7380 0.5317 0.7014 0.5030 0.5680 0.6074
Gemini 1.5 Flash 1 0.5753 0.7668 0.5272 0.6646 0.4548 0.5369 0.5876
3 0.5869 0.7686 0.5337 0.6789 0.4705 0.5473 0.5977
5 0.5868 0.7690 0.5303 0.6734 0.4661 0.5498 0.5959
10 0.5852 0.7665 0.5279 0.6773 0.4509 0.5464 0.5924
Table 3: Accuracy scores across six regional datasets for the ablation study (Section 5.3.2). Num($K$) ($k\in\{1,3,5,10\}$) indicates the number of demographic summaries retrieved. Bold text indicates the best performance, underlined text indicates the second-best performance.
@@ 5.1 Setup
Models Used.
We utilize GPT-4o-mini Achiam et al. (2023) and Gemini-1.5-Flash Team et al. (2024) for our generation tasks, which are accessed via APIs. We set the temperature parameter of these models to $0.7$ to achieve a balance between coherence and creativity. For the retrieval task, we employed the E5 (base) model Wang et al. (2022) [^4] https://huggingface.co/intfloat/e5-base, which generates embeddings and retrieves the top 100 most relevant summaries of values based on cosine similarity. Additionally, we utilize the GTE-multilingual-reranker-base model Zhang et al. (2024) [^5] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base for reranking. This reranking model refines the selection process and ensures that the top-$k$ most relevant summaries are selected from the initial top-100 retrievals.
Baseline Methods and Implementation.
Our baseline methods include: (1) Zero-shot inference, (2) the role-assignment-only approach Tao et al. (2024), (3) a few-shot learning method Choenni and Shutova (2024), and (4) a hybrid method that combines both (1) and (2).
Specifically, for the role-assignment baseline, we use the same demographic summaries as in ValuesRAG to ensure fairness by assigning roles based on demographic information from the survey data. For the few-shot method, we follow the approach outlined in the previous work, where we randomly select five examples from the test set as prompts. The hybrid method combines both strategies, assigning roles based on demographic summaries and augmenting the prompts with five randomly selected few-shot examples from the test set. [^6] Detailed prompts are provided in Appendix B.
Additionally, We use ValuesRAG with $k=3$ retrieved summaries—chosen to provide a good balance between retrieval diversity and contextual relevance. [^7] Detailed analysis of varying $k$ and its implications on model performance is provided in Section 5.3.1.
Evaluation.
We utilize accuracy as the primary evaluation metric by converting multiple-choice responses into a binary format for consistency and simplicity. This is achieved by categorizing responses into two distinct groups based on contrasting answer patterns, effectively capturing agreement and disagreement. This approach aligns with the structure of values-related questions, such as those using a Likert Scale.
@@ 5.2 Experimental Analysis
In our experiments, we compare our method, ValuesRAG, with four baseline methods: zero-shot inference, role-assignment, few-shot learning, and a hybrid approach that combines role-assignment and few-shot learning, as shown in Table 2.
We find that the role-assignment method generally surpasses both zero-shot and few-shot approaches. By grounding the agent’s responses in a clearly defined demographic context, it ensures more consistent performance. Yet, role assignment can sometimes lead to overly narrow representations when demograhpic roles are interpreted stereotypically. Meanwhile, few-shot learning can incorporate example-driven context, but its limited number of prompts may not consistently address the intricate ways individuals’ belifs diverge within similar social settings. As a result, it struggles to generalize to the multifaceted nature of human values, particularly when faced with unexpected or complex cultural scenarios. The hybrid method, which merges role assignment and few-shot prompts, does offer a partial improvement in contextual diversity. However, it remains insufficient for capturing the full spectrum of nuances that can arise from overlapping demographic factors and idiosyncratic personal persepectives.
@@
In contrast, ValuesRAG overcomes these challenges by dynamically retrieving and integrating specific values-related cultural data for each agent. By focusing on values as the primary retrieval targets, this retrieval-augmented framework enables the model to include an expansive set of contextual clues, helping it reflect the depth and breadth of each individual’s background and values. Crucially, our ValuesRAG provides a more adaptive mechanism for representing the subtle interplay of personal beliefs and cultural norms via avoiding the limitations of rigid demographic labels or small-sample prompts. ValuesRAG more effectively captures the complex dynamics that can shape a respondent’s stance on different questions. Evaluations across diverse test datasets demonstrate that ValuesRAG with $k=3$ consistently outperforms baseline methods, highlighting its ability to better represent cultural diversity, improve contextual alignment, and enhance overall model performance.
5.3 Ablation Study
We conduct two ablation studies to analyze the configuration and robustness of ValuesRAG., in Section 5.3.1, we vary the number of retrieved summaries ($k$) to examine how retrieval depth affects the model’s performance. Second, in Section 5.3.2, we isolate the effect of using only values-based generation.
5.3.1 5.3.1 Varying the Number of Retrieved Summaries
To further investigate the role of retrieval depth in ValuesRAG, we vary the number of retrieved summaries, $k\in\{1,3,5,10\}$. Table 3 presents the performance of ValuesRAG for three models across multiple datasets. We observe that while $k=1$ may not provide enough diversity, increasing $k$ beyond 3 can sometimes introduce less relevant information.
Moreover, $k=3$ consistently achieves significantly better accuracy compared to higher values of $k$, highlighting its effectiveness and achieving the best balance between contextual relevance and model latency; these ensure both effectiveness and efficiency. This configuration not only enhances retrieval accuracy but also maintains low computational overhead, underscoring the strength of ValuesRAG in real-world applications where speed and performance are equally crucial. Hence, we choose $k=3$ as the default retrieval number for conducting our baseline experiments, ensuring optimal performance across various models and datasets.
5.3.2 5.3.2 Impact of Values-Only Generation
@@ 5.3 Ablation Study
To validate the robustness of ValuesRAG, we perform an ablation study using only values context augmented generation, thereby excluding the impact of demographic summaries on the model’s performance. We use the WVS validation set—separated from the training data, which served as the retrieval corpus (as outlined in Section 3)—to evaluate the models. Table 4 presents a comparison of our method, using only summaries of values, against four baseline methods. Notably, ValuesRAG consistently outperforms all baselines across this validation data, achieving the highest accuracy despite relying exclusively on the values summaries.
Methods GPT-4o-mini Gemini-1.5-flash
Zero-Shot 0.6176 0.6041
Role-Assignment 0.6747 0.6505
Few-Shot Learning 0.6359 0.6086
Hybrid Method 0.6670 0.6354
Values Augmented 0.6894 0.6583
@@
This result confirms the effectiveness and robustness of the values-augmented generation approach. ValuesRAG leverages structured values summaries to generate contextually rich and culturally aligned responses. Even without demographic augmentation, ValuesRAG achieves superior performance by dynamically capturing the underlying value patterns, demonstrating its ability to generalize across diverse cultural contexts without requiring predefined QA examples or demographic anchors. The results demonstrate the framework’s scalability and adaptability, effectively mitigating biases and generating culturally coherent outputs with minimal dependence on external context.
@@ 6 Conclusion
We present ValuesRAG, an innovative Retrieval-Augmented Generation framework that revolutionizes cultural values alignment through its comprehensive approach to values context analysis. While existing approaches rely on pre-trained knowledge with fixed demographic labels, often leading to oversimplified generalizations and biases, or few-shot prompting techniques that struggle with generalization due to limited correlations across value dimensions, ValuesRAG overcomes these weaknesses by dynamically retrieving and integrating granular demographic data with detailed individual-level cultural profiles. Our extensive evaluations demonstrate that ValuesRAG achieves an accuracy improvement of up to 21% over state-of-the-art baselines. By combining adaptive retrieval mechanisms with in-context learning and reranking strategies, ValuesRAG captures complex cultural nuances, reduces biases, and ensures contextually aligned responses. This structured design enables ValuesRAG to bridge the gap between generic LLM capabilities and the demands of culturally sensitive applications, providing a scalable and robust solution for real-world use cases.
@@ 7 Limitation
Appendix A Appendix A. Dataset Description
A.1 WVS
As shown in Figure 3, the WVS dataset conducts surveys in over 120 countries worldwide, covering most regions and countries across the globe, making it suitable for constructing a documents corpus for Retrieval Augmented Generation (RAG). It is also representative in terms of demographic characteristics. We select the most recent WVS survey (Wave 7) to ensure the dataset is as close as possible to the contemporary world.
Figure 3: Distribution of Country in WVS dataset. The WVS dataset is the most extensive values survey dataset in terms of coverage.
According to the classification in the codebook released by WVS, the values-related questions it contains can be divided into 13 topics, as shown in Table 5, covering most dimensions of values, allowing for a comprehensive and accurate measure of each respondent’s values.
Topic Count
Social Values, Norms, Stereotypes 45
Happiness and Wellbeing 11
Social Capital, Trust and Organizational Membership 47
Economic Values 6
Perceptions of Corruption 9
Perceptions of Migration 10
Perceptions of Security 21
Index of Postmaterialism 6
Perceptions about Science and Technology 6
Religious Values 12
Ethical Values 23
Political Interest and Political Participation 35
Political Culture and Political Regimes 25
Table 5: Distribution of Values-related Questions in WVS. The questions were categorized into 13 topics with a total of 259 questions covering most of the dimensions of values
A.2 Test Datasets
To validate the robustness of our method, we selected datasets from six different regions as test sets. These test sets include both values-related questions and demographic characteristics. The demographic characteristics are used to generate summaries, serving as retrieval targets for RAG. The values-related questions are utilized as test questions to calculate accuracy (ACC).
EVS
The first dataset comes from the European Values Study [^8] https://europeanvaluesstudy.eu, a large-scale, cross-national, and longitudinal survey research program designed to explore values, beliefs, and attitudes across Europe. This dataset includes a total of 211 values-related questions and captures 34 demographic characteristics of the respondents. We select EVS 2017, conducted in 2017, ensuring alignment with the World Values Survey (WVS) in terms of the timeframe.
GSS
We select GSS to represent the population of the United States. The General Social Survey [^9] https://gss.norc.org is a sociological survey that has been conducted since 1972 by the National Opinion Research Center (NORC) at the University of Chicago. Its primary purpose is to collect and analyze data on the opinions, behaviors, and demographic characteristics of adults in the United States, thereby monitoring societal change and the growing complexity of American society. Its questionnaire covers a comprehensive and wide range of topics, including many values-related questions. Specifically, within the GSS, we identify 44 questions as values-related and 33 questions as demographic characteristics.
CGSS
The Chinese General Social Survey [^10] http://cgss.ruc.edu.cn, initiated in 2003, is China’s earliest national, comprehensive, and continuous academic survey project. Conducted by the National Survey Research Center at Renmin University of China, the CGSS systematically collects data at multiple levels, including society, communities, families, and individuals. CGSS only provided the questionnaire and data in Chinese, which we have translated into English to ensure its usability. We ultimately compile a total of 58 values-related questions and 13 demographic characteristics.
ISD
To ensure that our experiment covers as much of the world’s population as possible, we made efforts to include India within the scope of our test set. However, we were unable to obtain data from several government surveys in India, so we used data published by the Pew Research Center instead. The Pew Research Center’s India Survey Dataset [^11] https://www.pewresearch.org/dataset/india-survey-dataset is a comprehensive resource that captures the perspectives of 29,999 Indian adults on various aspects of society, including religious beliefs and practices, identity, nationalism, and societal tolerance. Conducted through face-to-face interviews between November 17, 2019, and March 23, 2020, the survey encompassed participants from diverse religious backgrounds, such as Hindus, Muslims, Sikhs, Christians, Buddhists, Jains, and others. This dataset covers 33 values-related questions and 23 demographic characteristics.
LAPOP
The Latin American Public Opinion Project (LAPOP) [^12] https://www.vanderbilt.edu/lapop is a research institute based at Vanderbilt University in Nashville, Tennessee. LAPOP’s most notable survey is the AmericasBarometer, the most extensive survey of democratic public opinion and behavior covering the Americas, including North, Central, South America, and the Caribbean. This survey measures democratic values and behaviors through voter surveys, providing valuable insights into public sentiments across the region. We select this dataset to represent the population of Latin America. There are 48 values-related questions and 12 demographic characteristics.
Africa
Afrobarometer [^13] https://www.afrobarometer.org is a pan-African, non-partisan research network established in 1999 that conducts public attitude surveys on democracy, governance, economic conditions, and related issues across Africa. We selected data from the 8th round of Afrobarometer (collected in 2022), covering 34 African countries. After screening, a total of 144 values-related questions and 14 demographic characteristics are obtained.
Appendix B Appendix B. Prompts Used
This section provides the prompts designed for various components of the ValuesRAG, including prompts for performing question answering tasks, as well as for generating values and demographic summaries.
Prompt for Question Answering Task: Respond to the question as the target individual, selecting the answer that aligns with their values and demographic context. Rules: Step-by-step analysis using retrieved demographics and values data. Maintain the target individual’s perspective throughout the analysis. Provide the response in JSON format, without additional explanation. Steps for Inferring: 1. Analyze the demographics (age, gender, cultural background, social class, religion, and economic class) of retrieved individuals. Compare them with the target individual. 2. Identify individuals whose demographics most closely match the target individual. Note their IDs. 3. Based on the matched individual’s values, infer how the target would respond. 4. Select the response that best aligns with the inferred values, and return only the integer representing the selected option.
Prompt for Values Summary Generation You are a summarization expert with expertise in extracting key insights from complex data. Based on the provided context, summarize this person’s values in one paragraph.
Prompt for Demographic Summary Generation You are a summarization expert with expertise in extracting key insights from complex data. Based on the provided context, summarize this person’s demographics in one paragraph.
Appendix C Appendix C. Case Study
Example 1 Question Is premarital sex always wrong, almost always wrong, sometimes wrong, or not wrong at all? Original Demographics Summary The respondent is a 63-year-old male born in the United States, of White descent, who speaks English at home. He completed 12th grade and holds a high school diploma, while his spouse has the same educational background. They are currently married and live in a household of two, with no children. The respondent identifies as middle class and reported a family income of \$90,000 to \$109,999 last year, with his own earnings falling within the same range. He practices Protestantism and is affiliated with the United Presbyterian Church in the U.S.A. His mother has completed two years of college, while his father finished high school. The respondent’s demographic background reflects a stable family life with no financial challenges reported. Retrieved Summaries: Retrieved Text 1
Demographics
The respondent is a 60-year-old White, non-Hispanic male born in the United States, where he speaks English at home. He was raised in a family where both parents were also born in the U.S., and he holds a bachelor’s degree, signaling a level of educational attainment typical of the upper middle class. He is currently employed full-time in a professional and technical role within the private business sector and is the chief wage earner in his household, which consists of one person. The respondent is divorced and has two children. He identifies as a Protestant and has reported saving money over the past year, rating his household income as relatively high on a scale of one to ten.
Values
This person’s values center on family, personal empowerment, and social equity, emphasizing a balanced life encompassing work, political engagement, and community well-being. They prioritize financial responsibility, health, and individual liberties while advocating equality and rejecting traditional gender roles. They approach trust cautiously, emphasize accountability in governance, and hold nuanced views on immigration and moral integrity. They believe in science and technology’s benefits and maintain personal faith amid diverse beliefs. Politically active, they support democratic principles and civil rights while voicing dissatisfaction with current institutions. Overall, they blend progressive values focused on social justice, environmental sustainability, and the power of citizen engagement in politics and public life.
Retrieved Text 2
Demographics
The respondent is a 63-year-old White, non-Hispanic male born in the United States, where he speaks English at home. He is a citizen of the U.S., with both parents also being born in the country. He is married and lives in a household of two people, including himself and his spouse, who is retired. He has two children and both he and his spouse completed 12th grade, earning a high school diploma. He works full-time in a semi-skilled occupation within the private sector and identifies as belonging to the middle class, reporting that his household income places them in the third income tier. The respondent aligns with the Protestant religion and describes their financial situation as stable.
Values
This individual values family and work as life’s center, while moderately engaging in politics. They emphasize independence, hard work, tolerance, and determination in children, embrace progressive stances on gender and social issues, and place little importance on religious conformity. Their trust is selective, relying heavily on family, showing moderate faith in others, and opposing corruption, while valuing economic equality and environmental concerns. They uphold science and technology, favor personal freedom over security, and maintain moral integrity and accountability. Strongly engaged in civic matters, they are committed to democracy, civil rights, and active political participation, yet remain cautious about government roles. Overall, their values balance familial dedication, personal agency, ethical living, and a progressive yet pragmatic outlook on social and political issues.
Retrieved Text 3
Demographics
The respondent is a 65-year-old white male, born in the United States, …
Example 2 Question How many children to have is a personal matter, and the government should not interfere. Do you agree? Original Demographics Summary The respondent is a 27-year-old male from China of Han ethnicity, who speaks Chinese at home. He holds a postgraduate degree and has never engaged in any work for economic income. Currently, he lives alone without any additional household members and identifies as belonging to the middle class, describing his household’s economic status as average in his local area. He does not practice any religion and reports no personal income in the past year. Retrieved Summaries: Retrieved Text 1
Demographics
The respondent is a 28-year-old male of Chinese ethnicity, born in the country where he currently resides. He speaks Standard Chinese (Mandarin/Putonghua/Guoyu) at home and identifies as part of the working class. He has attained a Bachelor’s degree and is employed full-time in a clerical role within the private sector. He lives in a household of four, including his own parents, and is currently married but has no children. Both he and his spouse have a similar educational background and are employed full-time. His family’s financial situation indicates that they have spent some savings and borrowed money over the past year. He does not belong to any religious denomination and has reported his household income as being in the third income level out of ten.
Values
This individual holds a complex set of values that place a significant emphasis on family, work, and personal responsibility, while also promoting income equality and valuing hard work as a means of success. They have a moderate stance on political engagement, expressing skepticism toward certain institutions while maintaining high trust in organizations like the police and government. Their views reflect a critical perspective on issues such as immigration and gender roles, where they value cultural diversity but prioritize native citizens for job opportunities. They exhibit a commitment to environmental protection, scientific advancements, and ethical living, while acknowledging their non-religious identity. Overall, their values convey a blend of personal autonomy, accountability, and a pragmatic view on governance, demonstrating a preference for democratic principles intertwined with the belief in the importance of social order and national security.
Retrieved Text 2
Demographics
The respondent is a 29-year-old male of Chinese descent, and speaks other Chinese dialects at home. He has attained a bachelor’s degree and works full-time in the sales sector within a private business or industry. Living alone, he is single and has no children. His family background includes parents with upper secondary education. Socially, he identifies with the working class and describes his household’s financial status as just getting by, placing his household income in the fourth income group on a scale of ten. He does not identify with any religious denomination.
Values
This person exhibits a set of values that emphasize the importance of family, hard work, and personal responsibility, reflecting a traditional view of societal roles and a commitment to instilling these qualities in children. They hold a pragmatic and somewhat secular worldview, expressing low interest in politics and religion while maintaining a positive attitude towards health, personal agency, and economic stability. They advocate for income equality and environmental sustainability, demonstrating concern for social equity and opposing corruption, while also valuing trust primarily within familial ties. Additionally, their perspective on immigrants is cautiously appreciative, and they prioritize freedom and economic progress, alongside a recognition of the complexities of modern ethical and moral issues. Overall, this individual’s values highlight a balance between familial loyalty, economic and environmental responsibilities, personal freedom, and a discerning approach to broader societal structures.
Retrieved Text 3
Demographics
The respondent is a 25-year-old male of Chinese ethnicity, who speaks Standard Chinese languages, …